Fermi National Laboratory


Accelerator Update

Monday July 19
The day shift began with Operations monitoring store 3661, with the Antiproton Source (Pbar) stacking, with Operations monitoring the Recycler stash, and with MiniBooNE taking beam.

At 8:36 AM, Booster RF technicians took down the east gallery to take BRF2 & 5 out of the RF system. Booster was back on line at 9:23 AM.

Pbar resumed stacking at 9:24 AM.

Operations established beam first to the Switchyard dump and then to the Meson MTest experiment at 10:44 AM.

RF techs brought a Main Injector RF station (MIRF12) on line at 11:28 AM.

At 1:04 PM, Booster RF techs took down the RF system to repairs BRF2. An hour later they had completed their work, tied the station back into the RF system, and Pbar resumed stacking.

Pbar experts began a study period at 2:30 PM. At 3:30 PM, experts began a reverse proton study.

At 4:26 PM, Operations reported that BRF1 had tripped off three times. About an hour later the station did this again, and then at 7 PM Operations reported that it happened several more times.

Pbar experts finished their studies at 9:12 PM.

At 9:15 PM, BRF2 stopped putting out RF so operators turned it off. Operators reconfigured the Booster RF system to work without it.

MTest experimenters said they were through using beam at 11:26 PM.

Tuesday July 20
The midnight shift began with Operations monitoring store 3661 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

At 1:22 AM, a Linac RF station (LRF5) tripped. It reset okay.

At 2 AM, another Linac RF station (KRF5) tripped of on a spark indication. It reset okay.

At 3:15 AM, KRF5 tripped of again on a spark indication. It reset okay.

At 3:49 AM, a Linac quadrupole power supply (L:Q2BULK) tripped off. A 15-volt control power supply failed. Operators replaced the power supply and had beam restored by 4:37 AM.

At 4:38 AM, an Accumulator SEPTA (A:ISEPV) tripped off. It reset okay.

Tevatron (TeV) experts began an end-of-store study at 5:05 AM. Unfortunately, the studies induced a TeV sector E1 quench at 5:16 AM.

At 6:09 AM, the MiniBooNE Horn tripped off for the first of three times in two hours. It reset okay each time. Operations reported that the trips were all Time Line Generator (TLG) related.

MCenter experimenters reported at 7:54 AM that they were finished with beam for the day shift.

At 8:23 AM, during a Mixed Mode - Antiproton shot setup, a MI RF station (MIRF12) began tripping off. RF techs looked at the station and made a quick update to the spark detector.

At 8:34 AM, LRF2 tripped off with no trouble indication. Operations reset the station.

At 8:43 AM, operators noticed that a quadrupole power supply current (QPS133) began jumping around when LRF2 tripped off. Linac experts investigated and adjusted LRF2's filament current.

Operations loaded the last protons into the TeV, after struggling with the Linac quadrupoles at 9:25 AM. Booster RF techs then took the RF system down to repair BRF2 at 9:48 AM.

At 10:04 AM, LRF2 went down for the first of three times in four minutes. It reset itself each time.

Operations established store 3663 at 10:10 AM with an initial luminosity of 91.55E30. Pbar resumed stacking.

A power supply tech began work on a Meson supply (MT5E) at 10:22 AM.

Operations began stashing antiprotons to Recycler at 11:28 AM.

Booster RF techs had BRF2 back online by 2:02 PM.

At 1:51 PM, A:ISEPV tripped off many times over a ten-minute period.

MI experts began slipstacking studies at 2:03 PM. At 2:12 PM, experts sent these high-intensity slipstacking pulses of beam to the Pbar target. The study ended at 5:10 PM.

Pbar resumed stacking at 5:50 PM.

MI suffered a vacuum burst at 7:38 PM. Operations reported at 7:58 PM that they were still trying to get the vacuum pumps to stay on.

At 7:47 PM, a Booster power supply (GMPS) tripped off and wouldn't reset. Experts came in and accessed the tunnel at 8:51 PM. They discovered and bypassed a faulty Klixon.

At 8:54 PM, Operations reported that they were having problems with the vacuum in the MI-52 area. Operators contacted an expert who came in to investigate.

At 10:15 PM, the Tevatron Electron Lens (TEL) tripped off. It reset okay.

At 10:51 PM, techs had to switch pumps for the Booster 95 LCW system. Operations turned off the Booster RF system for this switch, which was good because a number of stations tripped off. They all reset okay. The TEL, which operates off of the same water system, also tripped off at the same time. It reset okay.

At 11 PM, just as vacuum techs were about to access MI due to its vacuum problems, the techs managed to get the vacuum pumps to turn on and stay on. There was no need to access. This meant that the Recycler was able to keep its stash.

Pbar resumed stacking at 11:50 PM.

Wednesday July 21
The midnight shift began with Operations monitoring store 3663 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

Operations began preparing to send beam to the MCenter experiment at 1:05 AM.

At 1:50 AM, BRF11 tripped off. It reset okay, but then BRF6 tripped off. It reset okay.

At 5:42 AM, a TeV RF station (TRF8) tripped off.

At 6:03 AM, the I- Source tripped off due to no power. Operators investigated and found a problem with the motor generator. Operations hooked the dome to AC power to keep the temperatures stable and called in experts.

Plans for Wednesday Morning
The plans for today are to continue with stacking and storing. Operations will terminate the store early this evening and establish a Mixed Mode - Antiproton store. There will be a four-hour access beginning tomorrow morning around 8 AM for maintenance and repair. The store will remain in the TeV.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/21/2004   email Fermilab